# LapSum – One Method to Differentiate Them All: Ranking, Sorting and Top-k Selection

### v1

@@ 5.2 Experiments on soft-permutation methods
Baseline methods employ a common CNN to encode images into a feature space. The row-stochastic baseline combine CNN features into a vector processed by a multilayer perceptron for multiclass predictions. Sinkhorn and Gumbel-Sinkhorn approaches use the Sinkhorn operator to create doubly-stochastic matrices from these features. NeuralSort-based methods apply the NeuralSort operator with Gumbel noise to produce unimodal row-stochastic matrices. The loss in all methods is the row-wise cross-entropy against ground-truth permutation matrices with row-stochastic outputs. In our method $\mathrm{LapSum}$ we apply the same backbone architecture while eliminating the need for additional layers [^3] In $\mathrm{LapSum}$ model, logits are fed into a modified cross-entropy function to align the CNN’s output vector with the ground-truth permutation vector, reducing complexity while maintaining competitive performance. to transform the stochastic matrix. Tab. 4 provides results of this approach against other methods.
@@ Acknowledgments
This research was partially funded by the National Science Centre, Poland, grants no. 2023/49/B/ST6/01137 (work by Łukasz Struski and Jacek Tabor). Some experiments were performed on servers purchased with funds from the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the DigiWorld Priority Research Area within the Excellence Initiative – Research University program at Jagiellonian University in Kraków.
@@ Appendix A $\mathrm{LapSum}$: pseudocode for solution of top-k problem
$\sum_{i}p_{i}=w$
and the operation $s\to p$ is differentiable. In the limiting case when the pairwise differences between elements of the sequence $s_{i}$ are large, the solution tends to hard solution to top-w problem.
@@
Algorithm 1 LapSum:
soft top-$w$ function for the input $s=(s_{i})_{i=0..n-1}$
1:Sequence $(s_{i})_{i=0..n-1}$, parameter $w\in(0,n)$
@@
16:Find $k\in\{0,\dots,n\}$ such that $w\in[w_{k-1},w_{k}]$
17:if $k=0$ then
18:     $b=r_{0}-\log 2-\log w+\log a_{0}$
19:else if $0<k<n$ then
20:     $\begin{array}[]{l}b=r_{k+1}+\log a_{k}\\ -\log\big{(}w\!-\!c_{k+1}\!+\!\sqrt{|w\!-\!c_{k+1}|^{2}\!+\!a_{k}b_{k+1}e^{r_{k+1}-r_{k}}}\big{)}\end{array}$
@@ Appendix B Limiting case
For every $w\in(0,n)$ and $r\in\mathbb{R}^{n}$ we have
${F\text{-}\mathrm{Top}}_{\alpha}(r;w)\in\Delta_{w}.$ (6)
Moreover, if $k=w$ is integer and the elements of $r$ are pairwise distinct, then
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r;k)\to\mathop{\mathrm{topmin}}_{k}(r)$ as $\alpha\to 0^{+}$,
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r;k)\to\mathop{\mathrm{topmax}}_{k}(r)$ as $\alpha\to 0_{-}$.
@@
Observe that (6) follows directly from the fact that $b={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(w;r)$, and thus $\sum_{i}F_{\alpha}(b-r_{i})=w$.
We proceed to the proof of the limiting case, where we consider the case $\alpha\to 0^{+}$ (the case $\alpha\to 0^{-}$ can be shown analogously). Since the function ${F\text{-}\mathrm{Sum}}_{\alpha}$ is equivariant, without loss of generality we can assume that the sequence $(r_{i})$ is strictly increasing. Consider the values $b_{\alpha}={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(k;r)$. We show that
$F_{\alpha}(b_{\alpha}-r_{k})\to 0\text{ as }\alpha\to 0^{+}.$ (7)
If this would not be the case, then trivially there would exist a sequence $\alpha_{l}\to 0^{+}$ and $\varepsilon>0$ such that
$F_{\alpha_{l}}(b_{\alpha_{l}}-r_{k})=F(\frac{b_{\alpha_{l}}-r_{k}}{\alpha})\geq\varepsilon\text{ for all }l.$
By the continuity and fact that $F$ is strict increasing this implies that $\frac{b_{\alpha_{l}}-r_{k}}{\alpha}\geq c=F^{-1}(\varepsilon)$. Consequently for $i<k$, since $r_{k}-r_{i}>0$, we would get
$F_{\alpha_{l}}(b_{\alpha_{l}}-r_{i})$ $=F(\tfrac{b_{\alpha_{l}}-r_{i}}{\alpha_{l}})=F(\tfrac{b_{\alpha_{l}}-r_{k}}{\alpha_{l}}+\tfrac{r_{k}-r_{i}}{\alpha_{l}})$
$\geq F(c+\tfrac{r_{k}-r_{i}}{\alpha_{l}})\to 1\text{ as }l\to\infty$
@@
$\lim_{l\to\infty}\sum_{i=0}^{k-1}F_{\alpha_{l}}(b_{\alpha_{l}}-r_{i})+F_{\alpha_{l}}(b_{\alpha_{l}}-r_{k})\geq k+\varepsilon>k.$
and consequently to contradiction with the equality $\sum_{i}F_{\alpha}(b-r_{i})=k$.
Now this proves the assertion, since we trivially obtain that $F_{\alpha}(b-r_{j})\to 0$ for $\alpha\to 0^{+}$ and $j\geq k$. Consequently, making use of equality $\sum_{i}F_{\alpha}(b-r_{i})=k$ we obtain that $F_{\alpha}(b-r_{j})\to 1$ for $\alpha\to 0^{+}$ and $j<k$. ∎
@@ Appendix C Calculation of derivatives
In this section we are going to obtain main formulas for derivatives. We show in particular that all derivatives can be easily directly computed. Moreover, what is important for the gradient optimization process the multiplication of all the derivatives on a vector can be computed in $O(n\log n)$ time.
We start with establishing some notation. We are given $r=(r_{i})_{i=0..N-1}\in\mathbb{R}^{n}$, $w\in(0,N)$ and $\alpha\neq 0$. We consider the function
$b=b(r,w,\alpha)={Lap\text{-}\mathrm{Sum}}^{-1}(w;r,\alpha)$
@@
which for $\alpha>0$ corresponds to the density of the Laplace distribution.
@@
$s_{i}$ $=Lap_{\alpha}(b(r,w)-r_{i})=\tfrac{1}{2\alpha}\exp(-|\tfrac{x-r_{i}}{\alpha}|)$
@@
$=\mathrm{softmax}(-|\tfrac{x-r_{0}}{\alpha}|,\ldots,-|\tfrac{x-r_{N-1}}{\alpha}|).$
@@ C.1 Derivative of inverse of ${Lap\text{-}\mathrm{Sum}}$
Derivative of ${Lap\text{-}\mathrm{Sum}}^{-1}$ with respect to $w$
Differentiating (8) with respect to $w$, we get
$\sum_{j}Lap(b(r,w,\alpha)-r_{j})\cdot\frac{\partial b}{\partial w}=1,$
@@
$\frac{\partial b}{\partial w}=\frac{1}{\sum_{j}Lap_{\alpha}(b(w;r)-r_{j})}=\frac{1}{S}.$ (9)
@@
$\frac{\partial b}{\partial\alpha}$ $=\frac{1}{\alpha}\frac{\sum_{i}Lap\left(\frac{b-r_{i}}{\alpha}\right)(b-r_{i})}{\sum_{i}Lap\left(\frac{b-r_{i}}{\alpha}\right)}$
$=\frac{1}{\alpha}(b-\langle q,r\rangle).$
@@
$\sum_{j}Lap_{\alpha}(b(w;r))\cdot\left(\frac{\partial b}{\partial r_{i}}-1\right)=0,$
@@
$\frac{\partial b}{\partial r_{i}}\sum_{j}Lap_{\alpha}(b(w;r)-r_{j})=Lap_{\alpha}(b(w;r)-r_{i}),$
@@
$\frac{\partial b}{\partial r_{i}}=\frac{Lap_{\alpha}(b(w;r)-r_{i})}{\sum_{j}Lap_{\alpha}(b(w;r)-r_{j})}=q_{i},$
@@ C.2 Derivatives of ${Lap\text{-}\mathrm{Top}}$
$p_{i}(r,w,\alpha)=Lap_{\alpha}(b-r_{i})\text{ where }b={Lap\text{-}\mathrm{Sum}}^{-1}(r,w,\alpha).$
@@
Derivatives of $p_{i}(r_{1},\ldots,r_{n};w)$ with respect to $w$
@@
$\frac{\partial p_{i}}{\partial w}=\frac{\partial}{\partial w}Lap_{\alpha}(b(r;w)-r_{i})=Lap_{\alpha}(b(r;w)-r_{i})\cdot\frac{\partial b}{\partial w}.$
@@
$\frac{\partial p_{i}}{\partial w}=\frac{Lap_{\alpha}(b(r;w)-r_{i})}{\sum_{j}Lap_{\alpha}(x_{j}-b(x,w))}.$
@@
$\frac{dp}{dw}=q.$
Derivative of $p(\cdot;w)$ with respect to $\alpha$
@@
$\frac{\partial p_{i}}{\partial\alpha}$ $=\frac{\partial}{\partial\alpha}Lap_{\alpha}(b(r;w)-r_{i})=\frac{\partial}{\partial\alpha}Lap(\frac{b-r_{i}}{\alpha})$
$=Lap_{\alpha}(b(r;w)-r_{i})\cdot[\frac{\partial b}{\partial\alpha}-\frac{b-r_{i}}{\alpha}].$
@@
$\frac{\partial p_{i}}{\partial\alpha}$ $=Lap_{\alpha}(b(r;w)-r_{i})\cdot\left(\frac{b}{\alpha}-\frac{1}{\alpha}\langle q,r\rangle-\frac{b}{\alpha}+\frac{r_{i}}{\alpha}\right)$
@@
Derivative $D$ of $p(\cdot;w)$ with respect to $r$
@@
$d_{ij}=\frac{\partial p_{j}}{\partial r_{i}}=Lap_{\alpha}(b(w;r)-r_{i})\cdot\left(\frac{\partial b}{\partial r_{i}}-\delta_{ij}\right).$
@@
Derivative of $p(\cdot;w)$ in the direction $v$
@@ C.3 Computing derivatives for the function $u=\log p$
In some cases, especially for CE loss, we will need rather $\log p$ instead of $p$.
@@
$\log p_{i}=g(\frac{b-r_{i}}{\alpha}),$
@@
Derivative with respect to $w$
@@
$\frac{\partial u_{i}}{\partial w}=\frac{1}{p_{i}}\cdot\frac{\partial p_{i}}{\partial w}=\frac{q_{i}}{p_{i}}.$
@@
$\frac{\partial u_{i}}{\partial w}=q/p=h(p)/(S\alpha).$
@@
$Duv$ – Derivative with respect to $r$ in the direction $v$
@@
Refer to captionRefer to caption
(a) Results for CIFAR-100 with $P_{k}=[0,0,0,0,1]$ (top) and $P_{k}=[0.2,0.2,0.2,0.2,0.2]$ (bottom). The plots show Top-1 and Top-5 accuracy across varying values of $\alpha$.
Refer to captionRefer to caption
(b) Results for ImageNet-1K with $P_{k}=[0,0,0,0,1]$ (top) and $P_{k}=[0.5,0,0,0,0.5]$ (bottom). The plots show Top-1 and Top-5 accuracy across varying values of $\alpha$.
Figure 8: Top-1 and Top-5 accuracy of $\mathrm{LapSum}$ for CIFAR-100 (left) and ImageNet-1K (right) across varying $\alpha$ and $P_{k}$.

### v2

@@ 5.2 Experiments on soft-permutation methods
Baseline methods employ a common to all CNN to encode images into a feature space. The row-stochastic baseline combine CNN features into a vector processed by a multilayer perceptron for multiclass predictions. Sinkhorn and Gumbel-Sinkhorn approaches use the Sinkhorn operator to create doubly-stochastic matrices from these features. NeuralSort-based methods apply the NeuralSort operator with Gumbel noise to produce unimodal row-stochastic matrices. The loss in all methods is the row-wise cross-entropy against ground-truth permutation matrices with row-stochastic outputs. In our method $\mathrm{LapSum}$ we apply the same backbone architecture while eliminating the need for additional layers [^5] In $\mathrm{LapSum}$, logits are fed into a modified cross-entropy function to align the CNN’s output vector with the ground-truth permutation vector, reducing complexity while maintaining competitive performance. to transform the stochastic matrix. Tab. 4 provides results of this approach against other methods.
@@ Acknowledgments
This research was partially funded by the National Science Centre, Poland, grants no. 2020/39/D/ST6/01332 (work by Łukasz Struski and Michał B. Bednarczyk) and 2023/49/B/ST6/01137 (work by Jacek Tabor). Some experiments were performed on servers purchased with funds from the flagship project entitled “Artificial Intelligence Computing Center Core Facility” from the DigiWorld Priority Research Area within the Excellence Initiative – Research University program at Jagiellonian University in Kraków.
@@ Appendix A $\mathrm{LapSum}$: pseudocode for solution of top-k problem
$\sum_{i}p_{i}=k$
and the operation $s\to p$ is differentiable. In the limiting case when the pairwise differences between elements of the sequence $s_{i}$ are large, the solution tends to hard solution to top-k problem.
@@
Algorithm 1 LapSum: soft top-$k$ function
for the input $s=(s_{i})_{i=0..n-1}$
1:Sequence $(s_{i})_{i=0..n-1}$, parameter $\in(0,n)$
@@
16:Find $j\in\{0,\dots,n\}$ such that $k\in[w_{j-1},w_{j}]$
17:if $j=0$ then
18:  $b=r_{0}-\log 2-\log{}k+\log{}a_{0}$
19:else if $0<j<n$ then
20:   $\begin{array}[]{l}b=r_{j+1}+\log a_{j}\\ -\log\left(k\!-\!c_{j+1}\!+\!\sqrt{\left|k\!-\!c_{j+1}\right|^{2}\!+\!a_{j}b_{j+1}e^{r_{j+1}-r_{j}}}\right)\end{array}$
@@ Appendix B Limiting case
For every $k\in(0,n)$ and $r\in\mathbb{R}^{n}$ we have
${F\text{-}\mathrm{Top}}_{\alpha}(r,k)\in\Delta_{k}.$ (7)
Moreover, if $k$ is integer and the elements of $r$ are pairwise distinct, then
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r,k)\to\mathop{\mathrm{topmin}}_{k}(r)$ as $\alpha\to 0^{+}$,
  • ${F\text{-}\mathrm{Top}}_{\alpha}(r,k)\to\mathop{\mathrm{topmax}}_{k}(r)$ as $\alpha\to 0_{-}$.
@@
Observe that (7) follows directly from the fact that $b={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k)$, and thus $\sum_{i}F_{\alpha}(b-r_{i})=k$.
We proceed to the proof of the limiting case, where we consider the case $\alpha\to 0^{+}$ (the case $\alpha\to 0^{-}$ can be shown analogously). Since the function ${F\text{-}\mathrm{Sum}}_{\alpha}$ is equivariant, without loss of generality we can assume that the sequence $(r_{i})$ is strictly increasing. Consider the values $b_{\alpha}={F\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k)$. We show that
$F_{\alpha}(b_{\alpha}-r_{j})\to 0\text{ as }\alpha\to 0^{+}.$ (8)
If this were not the case then, trivially, there would exist a sequence $\alpha_{m}\to 0^{+}$ and $\varepsilon>0$ such that
$F_{\alpha_{m}}(b_{\alpha_{m}}-r_{j})=F\left(\frac{b_{\alpha_{m}}-r_{j}}{\alpha}\right)\geq\varepsilon\text{ for all }m.$
By continuity and fact that $F$ is strictly increasing, this implies that $\frac{b_{\alpha_{m}}-r_{j}}{\alpha}\geq c=F^{-1}(\varepsilon)$. Consequently for $i<k$, since $r_{j}-r_{i}>0$, we would get
$F_{\alpha_{m}}(b_{\alpha_{m}}-r_{i})$ $=F\left(\tfrac{b_{\alpha_{m}}-r_{i}}{\alpha_{m}}\right)=F\left(\tfrac{b_{\alpha_{m}}-r_{j}}{\alpha_{m}}+\tfrac{r_{j}-r_{i}}{\alpha_{m}}\right)$
$\geq F\left(c+\tfrac{r_{j}-r_{i}}{\alpha_{m}}\right)\to 1\text{ as }m\to\infty$
@@
$\lim_{m\to\infty}\sum_{i=0}^{k-1}F_{\alpha_{m}}(b_{\alpha_{m}}-r_{i})+F_{\alpha_{l}}(b_{\alpha_{m}}-r_{j})\geq k+\varepsilon>k.$
This we have obtained a contradiction with the equality $\sum_{i=0}^{n-1}F_{\alpha}(b-r_{i})=k$. This proves the assertion, since we trivially obtain that $F_{\alpha}(b-r_{j})\to 0$ for $\alpha\to 0^{+}$ and $j\geq k$. Consequently, making use of equality $\sum_{i}F_{\alpha}(b-r_{i})=k$ we obtain that $F_{\alpha}(b-r_{j})\to 1$ for $\alpha\to 0^{+}$ and $j<k$. ∎
@@ Appendix C Calculation of derivatives
In this section we are going to obtain main formulas for derivatives. We show, in particular, that all derivatives can be easily directly computed. Moreover, what is important for the gradient optimization process the multiplication of all the derivatives on a vector can be computed in $O(n\log n)$ time.
We shall start with establishing some consistent notation. We are given $r=(r_{i})_{i=0..n-1}\in\mathbb{R}^{n}$, $k\in(0,n)$ and $\alpha\neq 0$. We consider the function
$b=b(r,k,\alpha)={Lap\text{-}\mathrm{Sum}}^{-1}(r,k;\alpha)$
@@
which, for $\alpha>0$, corresponds to the density of the Laplace distribution.
@@
$s_{i}$ $=Lap_{\alpha}(b(r,k)-r_{i})=\tfrac{1}{2\alpha}\exp(-|\tfrac{x-r_{i}}{\alpha}|)$
@@
$=\mathrm{softmax}\left(-|\tfrac{x-r_{0}}{\alpha}|,\ldots,-|\tfrac{x-r_{n-1}}{\alpha}|\right).$
@@ C.1 Derivative of inverse of $Lap\text{-}\mathrm{Sum}$}
Derivative of ${Lap\text{-}\mathrm{Sum}}^{-1}$ with respect to $k$
Differentiating (9) with respect to $k$, we get
$\sum_{j}Lap_{\alpha}(b(r,w;\alpha)-r_{j})\cdot\frac{\partial b}{\partial w}=1,$
@@
$\frac{\partial b}{\partial k}=\frac{1}{\sum_{j}Lap_{\alpha}(b(r,k)-r_{j})}=\frac{1}{S}.$ (10)
@@
$\frac{\partial b}{\partial\alpha}$ $=\frac{1}{\alpha}\frac{\sum_{i}Lap_{\alpha}\left(\frac{b-r_{i}}{\alpha}\right)(b-r_{i})}{\sum_{i}Lap_{\alpha}\left(\frac{b-r_{i}}{\alpha}\right)}$
$=\frac{1}{\alpha}\left(b-\langle q,r\rangle\right).$
@@
$\sum_{j}Lap_{\alpha}(b(r,k))\cdot\left(\frac{\partial b}{\partial r_{i}}-1\right)=0,$
@@
$\frac{\partial b}{\partial r_{i}}\sum_{j}Lap_{\alpha}(b(r,k)-r_{j})=Lap_{\alpha}(b(r,k)-r_{i}),$
@@
$\frac{\partial b}{\partial r_{i}}=\frac{Lap_{\alpha}(b(r,k)-r_{i})}{\sum_{j}Lap_{\alpha}(b(r,k)-r_{j})}=q_{i},$
@@ C.2 Derivatives of $Lap\text{-}\mathrm{Top}$}
$p_{i}(r,w;\alpha)=Lap_{\alpha}(b-r_{i})\text{ where }b={Lap\text{-}\mathrm{Sum}}^{-1}(r,k;\alpha).$
@@
Derivatives of $p_{i}(r_{1},\ldots,r_{n},k)$ with respect to $k$
@@
$\frac{\partial p_{i}}{\partial k}=\frac{\partial}{\partial k}Lap_{\alpha}(b(r,k)-r_{i})=Lap_{\alpha}(b(r,k)-r_{i})\cdot\frac{\partial b}{\partial k}.$
@@
$\frac{\partial p_{i}}{\partial w}=\frac{Lap_{\alpha}(b(r,k)-r_{i})}{\sum_{j}Lap_{\alpha}(x_{j}-b(x,k))}.$
@@
$\frac{dp}{dk}=q.$
Derivative of $p(\cdot,k)$ with respect to $\alpha$
@@
$\frac{\partial p_{i}}{\partial\alpha}$ $=\frac{\partial}{\partial\alpha}Lap_{\alpha}(b(r,k)-r_{i})=\frac{\partial}{\partial\alpha}Lap\left(\frac{b-r_{i}}{\alpha}\right)$
$=Lap_{\alpha}(b(r,k)-r_{i})\cdot\left[\frac{\partial b}{\partial\alpha}-\frac{b-r_{i}}{\alpha}\right].$
@@
$\frac{\partial p_{i}}{\partial\alpha}$ $=Lap_{\alpha}(b(r,k)-r_{i})\cdot\left(\frac{b}{\alpha}-\frac{1}{\alpha}\langle q,r\rangle-\frac{b}{\alpha}+\frac{r_{i}}{\alpha}\right)$
@@
Derivative $D$ of $p(\cdot,k)$ with respect to $r$
@@
$d_{ij}=\frac{\partial p_{j}}{\partial r_{i}}=Lap_{\alpha}(b(r,k)-r_{i})\cdot\left(\frac{\partial b}{\partial r_{i}}-\delta_{ij}\right).$
@@
Derivative of $p(\cdot,k)$ in the direction $v$
@@ C.3 Computing derivatives for the function $u=\log p$
In some cases, especially for CE loss, we will rather need $\log p$ instead of $p$.
@@
$\log p_{i}=g\left(\frac{b-r_{i}}{\alpha}\right),$
@@
Derivative with respect to $k$
@@
$\frac{\partial u_{i}}{\partial k}=\frac{1}{p_{i}}\cdot\frac{\partial p_{i}}{\partial k}=\frac{q_{i}}{p_{i}}.$
@@
$\frac{\partial u_{i}}{\partial k}=q/p=h(p)/(\alpha\,{S}).$
@@
DuvDuv – Derivative with respect to $r$ in the direction $v$
@@
C.4 Derivatives of $(k_{i})_{i=0..L-1}\to({Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k_{i}))_{i=0..L-1}$
We shall now consider the case when the input is $k=(k_{i})\in\mathbb{R}^{L}$. This happens in particular in $\mathrm{LapSum}$ for soft sorting or the computation of soft permutation matrix.
We put
$b_{i}=b_{i}(r,k_{i};\alpha)={Lap\text{-}\mathrm{Sum}}^{-1}_{\alpha}(r,k_{i})$
and
$B=B(r,k;\alpha)=(b_{i})_{i=0..K-1}\in\mathbb{R}^{L}.$
Our aim is to compute the derivatives of the function $B$ with respect to $w$, $r$ and $\alpha$. We apply here the formulas from Sec. C.1.
We first introduce matrix $Q$ which is defined as
$Q=\begin{bmatrix}q_{0}^{T}\\ \vdots\\ q_{L-1}^{T}\end{bmatrix}$,
where $q_{m}\in\mathbb{R}^{n}$ is the (column) vector defined by
$q_{m}=\mathrm{softmax}(-\left|\tfrac{b_{m}-r_{0}}{\alpha}\right|,\ldots,-\left|\tfrac{b_{m}-r_{N-1}}{\alpha}\right|)\in\mathbb{R}^{n}.$
Derivative over $\alpha$
We have
$\frac{\partial b}{\partial\alpha}=\frac{1}{\alpha}(b-Q\cdot r).$
Derivative over $w$
We have
$\frac{\partial B}{\partial k}=\mathrm{diag}((1/S_{m})_{m=0..L-1}),$
where
$S_{m}=\frac{1}{2\alpha}\sum_{i}\exp(-\left|\tfrac{b_{m}-r_{i}}{\alpha}\right|)$
Derivative over $r$
We have
$\frac{db}{dr}=Q.$
Fast computation
Our aim is to show how one can efficiently compute the left and right multiplication of the above derivatives by vectors. We present here the general idea for fast computation of all $(S_{l})$, the other calculations can be done by applying a similar approach. For clarity of presentation, we restrict to the case $\alpha=1$.
So assume that we want to efficiently compute the values
$D_{m}=\sum_{i=0}^{n-1}\exp(-\left|b_{l}-r_{i}\right|)$
for all $m$, that is, in a smaller complexity than $O(L\cdot n)$. As before, we assume that the sequences $r$ and $b$ are in an increasing order. Observe that
$D_{m}=A_{m}+B_{m},$
where
$A_{m}=\sum_{i:r_{i}\leq b_{m}}\exp(r_{i}-b_{m}),\,B_{m}=\sum_{i:r_{i}>b_{m}}\exp(b_{m}-r_{i}).$
Let $n_{m}=\min\{i\in\{0,\ldots,n-1\}:r_{i}>b_{m}\}$ (and $n$ if there is no $i$ satisfying this condition). Then since the sequence $r$ is increasing we obtain that
$A_{m}=\sum_{i=0}^{n_{m}-1}\exp(r_{i}-b_{)},\,B_{m}=\sum_{i=n_{m}}^{n-1}\exp(b_{m}-r_{i}).$
Clearly, since sequences $r$ and $b$ are increasing we obtain that
$n_{m}\leq n_{m+1}.$
Consequently
$A_{m+1}=\exp(b_{m}-b_{m+1})\cdot A_{m}+\sum_{i=n_{m}}^{n_{m+1}}\exp(r_{i}-b_{m+1})$
and
$B_{m}=\exp(b_{m}-b_{m+1})\cdot B_{m+1}+\sum_{i=n_{m}}^{n_{m+1}}\exp(b_{m}-r_{i}).$
Therefore, we can first compute $A_{0}$ and $B_{n-1}$, and compute the rest by the above iterative formulas. Finally, the complexity of the computations is $O(L+n)$ instead of $O(L\cdot n)$.
Figure 8: ImageNet-1K Top-1 and Top-5 accuracies of $\mathrm{LapSum}$ across varying values of $\alpha$ and $P_{j}$. Larger $\alpha$’s produce ”harder”, more discrete selections while smaller $\alpha$ lead to smoother outputs. This trade-off influences how strictly the model enforces a top-$k$ criterion